### v1

@@ Abstract
Quadrupedal locomotion using deep reinforcement learning (DRL) is an enabling technology for terrain negotiation, stable mobility, and real-world deployment, and is a key capability in legged robotics. Existing approaches either train blind agents with heavy domain randomization, which sacrifices foresight, or adopt cross-modal fusion architectures that have their own limitations, where recurrent neural networks struggle to model long-term dependencies due to vanishing gradients, hierarchical designs complicate optimization and propagate errors across levels, and Transformer-based fusion incurs quadratic memory and computational cost from self-attention. To address these challenges, we introduce LocoMamba, a vision-driven cross-modal DRL framework built on selective state-space models, specifically leveraging Mamba, that achieves near-linear-time sequence modeling, effectively captures long-range dependencies, and enables efficient training with longer sequences. First, we embed proprioceptive states with a multilayer perceptron and patchify depth images with a lightweight convolutional neural network, producing compact tokens that improve state representation. Second, stacked Mamba layers fuse these tokens via near-linear-time selective scanning, reducing latency and memory footprint, remaining robust to token length and image resolution, and providing an inductive bias that mitigates overfitting. Third, we train the policy end-to-end with Proximal Policy Optimization under terrain and appearance randomization and an obstacle-density curriculum, using a compact state-centric reward that balances progress, smoothness, and safety. We evaluate our method in challenging simulated environments with static and moving obstacles as well as uneven terrain. Compared with state-of-the-art baselines, our method achieves higher returns and success rates with fewer collisions, exhibits stronger generalization to unseen terrains and obstacle densities, and improves training efficiency by converging in fewer updates under the same compute budget.
@@
Effective proprioception and depth encoders: We introduce a compact input representation in which an MLP embeds proprioceptive states and a CNN patchifies depth images into spatial tokens tailored for Mamba-based fusion. This design provides immediate state estimates and look-ahead while reducing sensitivity to appearance variation, thereby improving computational efficiency and training stability.
Efficient cross-modal Mamba fusion backbone: We fuse encoded tokens from proprioception and depth using stacked Mamba layers via selective state-space scanning, achieving near-linear time and memory growth. The backbone supports long-horizon modeling, remains robust to token length and image resolution, and provides a regularizing inductive bias through input-gated, exponentially decaying dynamics.
Robust end-to-end RL training scheme: We train the policy with PPO using Mamba-fused cross-modal features, complemented by terrain and appearance randomization and an obstacle-density curriculum. A compact, state-centric reward balances progress, smoothness, efficiency, and safety, enabling stable learning and consistent performance.
Comprehensive evaluation: We conduct extensive experiments with static and moving obstacles and uneven terrain, and demonstrate consistent gains over state-of-the-art (SOTA) in terms of return, success rate, and collision rate, along with faster convergence under the same compute budget.
@@ 1 Introduction
Quadrupedal robots provide mobility in environments where wheeled platforms are ineffective, such as stairs, rubble, soft or deformable substrates, and cluttered indoor or outdoor settings, enabling applications in inspection, disaster response, agriculture, and planetary exploration (Fan et al., 2024). Robust locomotion control is therefore a foundational capability for practical quadrupedal systems, underpinning safe navigation, dependable mission execution, and reliable operation across diverse terrains and disturbance conditions (Carpentier and Wieber, 2021).
@@
Current approaches to quadrupedal locomotion control fall into two broad categories. Classical model-based pipelines depend on accurate dynamics and terrain models, explicit contact planning, and extensive parameter tuning, which constrain scalability and complicate deployment (Xin et al., 2021). Learning-based methods mitigate this burden and enable end-to-end policies that couple perception with control (Ha et al., 2025). Within this paradigm, deep reinforcement learning has emerged as the dominant approach because it optimizes task-driven closed-loop policies through interaction with the environment and produces robust, adaptive behaviors (Zhang et al., 2022).
@@
DRL methods have substantially advanced quadrupedal locomotion, enabling traversal of uneven terrain (Xie et al., 2021) and operation under challenging conditions such as mud, snow, and running water (Lee et al., 2020). However, most studies continue to train blind controllers that rely only on proprioception and achieve stability primarily through large-scale simulation and domain randomization. Despite their robustness, blind agents lack foresight because they receive no exteroceptive input, so they react only after contact and struggle to avoid obstacles proactively or to plan foot placement on irregular ground (Xie et al., 2022).
@@
On the other hand, cross-modal fusion with visual input has gained traction. Vision complements proprioception by providing look-ahead, enabling early perception of distant obstacles, anticipation of terrain changes before contact, and timely trajectory adjustments that reduce collisions and improve foot placement (Han et al., 2025). Nevertheless, current fusion architectures exhibit notable trade-offs. Recurrent models often suffer from vanishing gradients and limited capacity for long-horizon dependencies (Li et al., 2024). Hierarchical designs complicate optimization and propagate errors across levels (Jain et al., 2019). Transformer-based fusion, although expressive, incurs quadratic memory use and computational cost with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency (Yang et al., 2021).
@@
In this paper, we propose LocoMamba, a vision-driven end-to-end DRL framework that uses Mamba as a selective state-space model (SSM) backbone to enable efficient quadrupedal locomotion. A lightweight multilayer perception (MLP) embeds proprioceptive states to provide accurate estimates of the robot state for immediate reaction, while a compact convolutional neural network (CNN) patchifies depth images to supply look-ahead for negotiating uneven terrain and large obstacles. Stacked Mamba layers fuse tokens through selective state-space scanning with near-linear-time complexity, replacing quadratic self-attention with recurrent updates that lower latency and memory use. The streaming formulation accommodates variable-length inputs, enabling longer visual context and higher token resolution, while input-gated, exponentially decaying dynamics provide a regularizing inductive bias that mitigates overfitting. Policies are trained with Proximal Policy Optimization (PPO) under terrain and appearance randomization together with an obstacle-density curriculum, which broadens environmental coverage and progressively increases difficulty to stabilize on-policy learning. The reward is compact and state-centric, encouraging progress and accurate tracking, promoting smooth and efficient actuation through torque and acceleration regularization, and enforcing safety through collision, fall, and foot-clearance terms.
@@
    Effective proprioception and depth encoders: We introduce a compact input representation in which an MLP embeds proprioceptive states and a CNN patchifies depth images into spatial tokens tailored for Mamba-based fusion. This design provides immediate state estimates and look-ahead while reducing sensitivity to appearance variation, thereby improving computational efficiency and training stability.
@@
    Efficient cross-modal Mamba fusion backbone: We fuse encoded tokens from proprioception and depth using stacked Mamba layers via selective state-space scanning, achieving near-linear time and memory growth. The backbone supports long-horizon modeling, remains robust to token length and image resolution, and provides a regularizing inductive bias through input-gated, exponentially decaying dynamics.
@@
    Robust end-to-end RL training scheme: We train the policy with PPO using Mamba-fused cross-modal features, complemented by terrain and appearance randomization and an obstacle-density curriculum. A compact, state-centric reward balances progress, smoothness, efficiency, and safety, enabling stable learning and consistent performance.
@@
    Comprehensive evaluation: We conduct extensive experiments with static and moving obstacles and uneven terrain, and demonstrate consistent gains over state-of-the-art (SOTA) in terms of return, success rate, and collision rate, along with faster convergence under the same compute budget.
@@ 2.1 Learning-Based Quadrupedal Locomotion
Research on quadrupedal locomotion has evolved along two principal paradigms: rule-based control and learning-based approaches. Early rule-based controllers relied on template dynamics and heuristic regulation of balance, leg compliance, and foot placement, establishing fundamental principles for dynamic locomotion (Miura and Shimoyama, 1984; Liu and Ding, 2014; Habu et al., 2018; Bledt et al., 2018). Subsequent rule-based systems formalized locomotion as constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control with explicit contact constraints and task hierarchies (Grandia et al., 2019; Amatucci et al., 2024; Elobaid et al., 2025). For example, (Di Carlo et al., 2018) developed real-time MPC and task-space whole-body control on torque-controlled quadrupeds, demonstrating stable trotting, stair traversal, and robust disturbance rejection. (Ding et al., 2019) refined MPC cost design and contact modeling to improve tracking and compliance under high-frequency control budgets. (Carius et al., 2019) studied trajectory optimization and contact-aware planning for foothold selection and force allocation, improving reliability on structured terrain. However, these approaches require accurate models of the environment and system dynamics and substantial manual tuning, thereby limiting their applicability in complex and variable settings.
On the other hand, model-free RL can learn general policies for challenging conditions (Li et al., 2021; Margolis et al., 2024; Bussola et al., 2025; Ha et al., 2025). For example, (Hwangbo et al., 2019) trained torque-level policies with dynamics randomization and privileged learning to achieve fast, robust trotting on uneven terrain. (Tan et al., 2018) used large-scale domain randomization and dynamics perturbations to transfer simulation-trained policies to real quadrupeds with strong disturbance rejection. (Kumar et al., 2021) introduced rapid motor adaptation that augments a proprioceptive policy with an online adaptation module, enabling quick recovery under payload shifts and terrain changes. However, most approaches are proprioception-only and thus blind to exteroceptive cues, which limits foresight for obstacle negotiation and precise foothold planning. In this work, we fuse visual and proprioceptive inputs to obtain a richer state representation, enabling the policy to anticipate terrain changes and plan trajectories online while maintaining continuous motion.
@@ 2.2 Vision-Driven RL for Quadrupedal Locomotion
To extend RL beyond state-only inputs, a growing body of work leveraged visual observations for locomotion control (Han et al., 2025). For example, Yu et al. (2021) trained a vision-based controller that incorporated exteroceptive input directly into the RL loop and achieved traversal of uneven terrain and complex obstacles. Duan et al. (2024) learned a controller in simulation with heightmap observations and then trained a depth-to-heightmap predictor from depth and state histories, enabling vision-guided locomotion with transfer to hardware on challenging terrains. Complementing these efforts, Fahmi et al. (2022) proposed the ViTAL framework, which used vision for terrain-aware planning by decomposing decision making into foothold selection and pose adaptation, and improved safety and reliability on irregular surfaces. Hierarchical formulations were also explored; for example, Jain et al. (2019) employed hierarchical RL in which high-level visual policies guided low-level motor control with sensor inputs.
Despite these advances, common fusion strategies involved trade-offs. Simple concatenation underused spatial structure, and hierarchical decompositions increased optimization complexity and risked error propagation across levels (Singh et al., 2022). In this work, we avoid complex hierarchical designs and adopt a simple end-to-end schema that fuses proprioceptive and depth inputs within a single policy trained with PPO, enabling visual look-ahead while maintaining a streamlined learning pipeline.
@@ 2.3 State-Space Models and Cross-Modal RL
Cross-modal RL has attracted significant attention. Typical fusion backbones include recurrent neural networks (RNNs) and Transformers. For example, (Xiao et al., 2024) designed an egocentric visuomotor RL framework that fused egocentric depth and proprioception with long short-term memory (LSTM) units, enabling a quadruped to negotiate obstacles and traverse cluttered scenes in the real world. (Lai et al., 2024) proposed a gated recurrent unit (GRU)–based world-model formulation that fuses depth images and proprioceptive readings into a recurrent latent state for visual legged locomotion, improving control under partial observability. (Yang et al., 2021) introduced a cross-modal Transformer policy that ingests depth tokens and proprioceptive features end-to-end, achieving stronger terrain anticipation and sim-to-real transfer than blind baselines; subsequent work scaled such tokenized proprioception-vision Transformer policies to heterogeneous robot datasets to improve generalization. However, RNNs often suffer from vanishing gradients and limited long-horizon capacity, which complicates optimization in extended sequences. Transformer models offer strong expressivity but incur quadratic memory and computation with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency.
SSMs offer an alternative by updating a compact recurrent state per token and scaling near-linearly with sequence length. Building on the S4 model (Gu et al., 2021), which demonstrated efficient long-range dependency modeling on long-sequence benchmarks, (Gu and Dao, 2023) introduced Mamba , a selective SSM whose input-dependent parameters and hardware-aware parallel scan yield linear-time sequence modeling with favorable throughput and memory characteristics across modalities.
Mamba-style SSMs have also been explored for multimodal fusion. AlignMamba augments a multimodal Mamba backbone with token-level optimal-transport alignment and a global distributional alignment objective to improve cross-modal consistency (Li et al., 2025). COMO (Cross-Mamba Interaction with Offset-Guided Fusion) addresses sensor misalignment for multimodal object detection while retaining the efficiency benefits of selective scanning (Liu et al., 2026). AV-Mamba applies selective SSMs to audio-visual question answering and reports improvements over Transformer baselines (Huang et al., 2024). FusionMamba (Xie et al., 2024) and DepMamba (Ye et al., 2025) further demonstrate SSM-based gains in multimodal image fusion and audio-visual affect analysis. To the best of our knowledge, this is the first work to use cross-modal Mamba for quadrupedal locomotion.
@@ 3 Methodology
Figure 1 presents the overall architecture of LocoMamba. The pipeline comprises three components trained end to end. First, proprioceptive states and depth images are encoded into a compact tokenized latent space: an MLP maps the proprioceptive vector to a state token, and a lightweight CNN patchifies the depth image into spatial tokens. Second, the concatenated token stream is fused by stacked Mamba SSM layers via selective state-space scanning, which updates a compact recurrent state per token, yields near linear time and memory scaling, accommodates variable token counts and resolutions, and preserves long-horizon temporal context. Third, policy and value heads consume the fused features and are optimized with PPO under a compact state-centric reward that balances progress, smoothness and efficiency, and safety.

### v2

@@ Abstract
Applying deep reinforcement learning (DRL) to quadrupedal robots is promising for obstacle negotiation, terrain locomotion, and real-world deployment. Existing methods either train blind agents that sacrifice foresight or adopt cross-modal fusion architectures with limitations; even Transformer-based models still suffer from the quadratic cost of self-attention. To overcome these challenges, we propose LocoMamba, a vision-driven cross-modal DRL framework built on selective state-space models (Mamba). LocoMamba encodes proprioceptive and vision features into compact tokens, which are fused by stacked Mamba layers through near-linear-time selective scanning. This reduces latency and memory usage while preserving long-range dependencies. The policy is trained end-to-end with Proximal Policy Optimization under randomized terrain and obstacle-density curriculum, using a state-centric reward balancing task and safety. Evaluations in diverse static and dynamic obstacle settings show that LocoMamba significantly outperforms a Transformer baseline, achieving $48.9\%$ higher returns, $30.4\%$ longer distance, and $48.9\%$ fewer collisions on trained terrains, while exhibiting stronger generalization to unseen scenarios. Moreover, it converges faster with limited compute, enhancing training efficiency. Overall, LocoMamba improves locomotion performance and reduces training cost, supporting rapid iteration and deployment of quadruped learning algorithms. A repository is hosted at https://github.com/allen-quad-robot/locomamba, to be made available upon acceptance.
@@
Effective proprioception and depth encoders: A compact input representation is introduced in which an MLP embeds proprioceptive states and a CNN patchifies depth images into spatial tokens tailored for Mamba-based fusion. This design provides immediate state estimates and look-ahead while reducing sensitivity to appearance variation, thereby improving computational efficiency and training stability.
Efficient cross-modal Mamba fusion backbone: Encoded tokens from proprioception and depth are fused using stacked Mamba layers via selective state-space scanning, achieving near-linear time and memory scaling. The backbone supports long-horizon modeling, remains robust to token length and image resolution, and provides a regularizing inductive bias through input-gated, exponentially decaying dynamics.
Robust end-to-end RL training scheme: The policy is trained with PPO using Mamba-fused cross-modal features, complemented by terrain and appearance randomization and an obstacle-density curriculum. A compact, state-centric reward balances task-aligned progress, energy efficiency, and safety, enabling stable learning and consistent performance.
Comprehensive evaluation: Extensive experiments are conducted with static and moving obstacles and uneven terrain, and demonstrate consistent gains over state-of-the-art (SOTA) in terms of return, collision times, and distance moved, along with faster convergence under the same compute budget.
@@ 1 Introduction
Quadrupedal robots provide mobility in environments where wheeled platforms are ineffective, such as stairs, rubble, soft or deformable substrates, and cluttered indoor or outdoor settings, enabling applications in inspection, disaster response, agriculture, and planetary exploration [1]. Robust locomotion control is therefore a foundational capability for practical quadrupedal systems, underpinning safe navigation, dependable mission execution, and reliable operation across diverse terrains and disturbance conditions [2].
@@
Current approaches to quadrupedal locomotion control fall into two broad categories. Classical model-based pipelines depend on accurate dynamics and terrain models, explicit contact planning, and extensive parameter tuning, which constrain scalability and complicate deployment [3]. Learning-based methods mitigate this burden and enable end-to-end policies that couple perception with control [4]. Within this paradigm, deep reinforcement learning has emerged as the dominant approach because it optimizes task-driven closed-loop policies through interaction with the environment and produces robust, adaptive behaviors [5].
@@
DRL methods have substantially advanced quadrupedal locomotion, enabling traversal of uneven terrain [6] and operation under challenging conditions such as mud, snow, and running water [7]. However, most studies continue to train blind controllers that rely only on proprioception and achieve stability primarily through large-scale simulation and domain randomization. Despite their robustness, blind agents lack foresight because they receive no exteroceptive input, so they react only after contact and struggle to avoid obstacles proactively or to plan foot placement on irregular ground [8].
@@
On the other hand, cross-modal fusion with visual input has gained traction. Vision complements proprioception by providing look-ahead, enabling early perception of distant obstacles, anticipation of terrain changes before contact, and timely trajectory adjustments that reduce collisions and improve foot placement [9]. Nevertheless, current fusion architectures exhibit notable trade-offs. Recurrent models often suffer from vanishing gradients and limited capacity for long-horizon dependencies [10]. Hierarchical designs complicate optimization and propagate errors across levels [11]. Transformer-based fusion, although expressive, incurs quadratic memory use and computational cost with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency [12].
@@
In this paper, LocoMamba, a vision-driven end-to-end DRL framework that uses Mamba as a selective state-space model (SSM) backbone, is introduced to enable efficient quadrupedal locomotion. A lightweight multilayer perception (MLP) embeds proprioceptive states to provide accurate estimates of the robot state for immediate reaction, while a compact convolutional neural network (CNN) patchifies depth images to supply look-ahead for negotiating uneven terrain and large obstacles. Stacked Mamba layers fuse tokens through selective state-space scanning with near-linear time complexity, using recurrent updates whose cost scales with sequence length and thereby lowering latency and memory use compared with quadratic self-attention. The streaming formulation accommodates variable-length inputs, enabling longer visual context and higher token resolution, while input-gated, exponentially decaying dynamics provide a regularizing inductive bias that mitigates overfitting. Policies are trained with Proximal Policy Optimization (PPO) under terrain and appearance randomization together with an obstacle-density curriculum, which broadens environmental coverage and progressively increases difficulty to stabilize on-policy learning. The reward is compact and state-centric, encouraging task-aligned progress, promoting energy-efficient actuation, and enforcing safety via penalties for collisions and falls.
@@
    Effective proprioception and depth encoders: A compact input representation is proposed in which an MLP embeds proprioceptive states and a CNN patchifies depth images into spatial tokens tailored for Mamba-based fusion. This design provides immediate state estimates and foresight while reducing sensitivity to appearance variation, thereby improving computational efficiency and training stability.
@@
    Efficient cross-modal Mamba fusion backbone: The encoded tokens are fused from proprioception states and depth images using stacked Mamba layers via selective state-space scanning, achieving near-linear time and memory scaling. The backbone supports long-horizon modeling, remains robust to token length and image resolution, and provides a regularizing inductive bias through input-gated, exponentially decaying dynamics.
@@
    Robust end-to-end RL training scheme: The policy is trained with PPO using Mamba-fused cross-modal features, complemented by terrain and appearance randomization and an obstacle-density curriculum. A compact, state-centric reward balances task-aligned progress, energy efficiency, and safety, enabling stable learning and consistent performance.
@@
    Comprehensive evaluation: Extensive experiments are conducted with static and moving obstacles and uneven terrain, and demonstrate consistent gains over state-of-the-art (SOTA) in terms of return, collision times, and distance moved, along with faster convergence under the same compute budget.
@@ 2.1 Learning-Based Quadrupedal Locomotion
Research on quadrupedal locomotion has evolved along two principal paradigms: rule-based control and learning-based approaches. Early rule-based controllers relied on template dynamics and heuristic regulation of balance, leg compliance, and foot placement, establishing fundamental principles for dynamic locomotion [13, 14, 15, 16]. Subsequent rule-based systems formalized locomotion as constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control with explicit contact constraints and task hierarchies [17, 18, 19, 20]. For example, [21] developed real-time MPC and task-space whole-body control on torque-controlled quadrupeds, demonstrating stable trotting, stair traversal, and robust disturbance rejection. [22] refined MPC cost design and contact modeling to improve tracking and compliance under high-frequency control budgets. [23] studied trajectory optimization and contact-aware planning for foothold selection and force allocation, improving reliability on structured terrain. However, these approaches require accurate models of the environment and system dynamics and substantial manual tuning, thereby limiting their applicability in complex and variable settings.
On the other hand, model-free RL can learn general policies for challenging conditions [24, 25, 26, 4]. For example, [27] trained torque-level policies with dynamics randomization and privileged learning to achieve fast, robust trotting on uneven terrain. [28] used large-scale domain randomization and dynamics perturbations to transfer simulation-trained policies to real quadrupeds with strong disturbance rejection. [29] introduced rapid motor adaptation that augments a proprioceptive policy with an online adaptation module, enabling quick recovery under payload shifts and terrain changes. However, most approaches are proprioception-only and thus blind to exteroceptive cues, which limits foresight for obstacle negotiation and precise foothold planning. In this work, both the visual and proprioceptive inputs are fused to obtain a richer state representation, enabling the policy to anticipate terrain changes and plan trajectories online while maintaining continuous motion.
@@ 2.2 Vision-Driven RL for Quadrupedal Locomotion
To extend RL beyond state-only inputs, a growing body of work leveraged visual observations for locomotion control [9]. For example, Yu et al. [30] trained a vision-based controller that incorporated exteroceptive input directly into the RL loop and achieved traversal of uneven terrain and complex obstacles. Duan et al. [31] learned a controller in simulation with heightmap observations and then trained a depth-to-heightmap predictor from depth and state histories, enabling vision-guided locomotion with transfer to hardware on challenging terrains. Complementing these efforts, Fahmi et al. [32] proposed the ViTAL framework, which used vision for terrain-aware planning by decomposing decision making into foothold selection and pose adaptation, and improved safety and reliability on irregular surfaces. Hierarchical formulations were also explored; for example, Jain et al. [11] employed hierarchical RL in which high-level visual policies guided low-level motor control with sensor inputs.
Despite these advances, common fusion strategies involved trade-offs. Simple concatenation underused spatial structure, and hierarchical decompositions increased optimization complexity and risked error propagation across levels [33]. In this work, complex hierarchical designs are avoided and a simple end-to-end schema is adopted that fuses proprioceptive and depth inputs within a single policy trained with PPO, enabling visual look-ahead while maintaining a streamlined learning pipeline.
@@ 2.3 State-Space Models and Cross-Modal RL
Cross-modal RL has attracted significant attention. Typical fusion backbones include recurrent neural networks (RNNs) and Transformers. For example, [34] designed an egocentric visuomotor RL framework that fused egocentric depth and proprioception with long short-term memory (LSTM) units, enabling a quadruped to negotiate obstacles and traverse cluttered scenes in the real world. [35] proposed a gated recurrent unit (GRU)–based world-model formulation that fuses depth images and proprioceptive readings into a recurrent latent state for visual legged locomotion, improving control under partial observability. [12] introduced a cross-modal Transformer policy that ingests depth tokens and proprioceptive features end-to-end, achieving stronger terrain anticipation and sim-to-real transfer than blind baselines; subsequent work scaled such tokenized proprioception-vision Transformer policies to heterogeneous robot datasets to improve generalization. However, RNNs often suffer from vanishing gradients and limited long-horizon capacity, which complicates optimization in extended sequences. Transformer models offer strong expressivity but incur quadratic memory and computation with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency.
SSMs [36] offer an alternative by updating a compact recurrent state per token and scaling near-linearly with sequence length. Building on the S4 model [37], which demonstrated efficient long-range dependency modeling on long-sequence benchmarks, [38] introduced Mamba , a selective SSM whose input-dependent parameters and hardware-aware parallel scan yield linear-time sequence modeling with favorable throughput and memory characteristics across modalities.
Mamba-style SSMs have also been explored for multimodal fusion. AlignMamba augments a multimodal Mamba backbone with token-level optimal-transport alignment and a global distributional alignment objective to improve cross-modal consistency [39]. COMO (Cross-Mamba Interaction with Offset-Guided Fusion) addresses sensor misalignment for multimodal object detection while retaining the efficiency benefits of selective scanning [40]. AV-Mamba applies selective SSMs to audio-visual question answering and reports improvements over Transformer baselines [41]. FusionMamba [42] and DepMamba [43] further demonstrate SSM-based gains in multimodal image fusion and audio-visual affect analysis. To the best of our knowledge, this is the first work to use cross-modal Mamba for quadrupedal locomotion.
@@ 3 Methodology
Fig. 1 presents the overall architecture of LocoMamba. The pipeline comprises three components trained end to end. First, proprioceptive states and depth images are encoded into a compact tokenized latent space: an MLP maps the proprioceptive vector to a state token, and a lightweight CNN patchifies the depth image into spatial tokens. Second, the concatenated token stream is fused by stacked Mamba SSM layers via selective state-space scanning, which updates a compact recurrent state per token, yields near linear time and memory scaling, accommodates variable token counts and resolutions, and preserves long-horizon temporal context. Third, policy and value heads consume the fused features and are optimized with PPO [44] under a compact state-centric reward that balances task-aligned progress, energy efficiency, and safety.